40m
QIL
Cryo_Lab
CTN
SUS_Lab
TCS_Lab
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Thu Jun 11 11:48:43 2020, gautam, Update, VAC, Vac failure
|
Thu Jun 11 16:14:03 2020, gautam, Update, VAC, Vac failure - probable cause is serial comm glitch  
|
Thu Jun 11 17:35:34 2020, gautam, Update, VAC, Pumpspool UPS needs battery replacement
|
Fri Jun 12 01:23:32 2020, Koji, Update, VAC, Pumpspool UPS needs battery replacement
|
Fri Jun 12 17:32:40 2020, Koji, Update, VAC, Pumpspool UPS needs battery replacement
|
Fri Jun 12 19:02:52 2020, gautam, Update, VAC, Pumpspool UPS needs battery replacement
|
Fri Jun 12 19:23:56 2020, Koji, Update, VAC, Pumpspool UPS needs battery replacement
|
Fri Jun 12 19:33:31 2020, gautam, Update, VAC, Pumpspool UPS needs battery replacement
|
Tue Jun 16 13:35:03 2020, Jon, Update, VAC, Temporary vac fix / IFO usable again
|
Wed Jun 17 16:27:51 2020, gautam, Update, VAC, Questions/comments on vacuum
|
Thu Jun 18 11:00:24 2020, Jon, Update, VAC, Questions/comments on vacuum
|
Thu Jun 18 12:00:36 2020, gautam, Update, VAC, Questions/comments on vacuum
|
Thu Jun 18 14:13:03 2020, Jon, Update, VAC, Questions/comments on vacuum
|
Thu Jun 18 15:46:34 2020, gautam, Update, VAC, Questions/comments on vacuum
|
Fri Jun 19 07:40:49 2020, Jon, Update, VAC, Questions/comments on vacuum
|
Fri Jun 19 09:57:35 2020, gautam, Update, VAC, Questions/comments on vacuum
|
|
Message ID: 15392
Entry time: Thu Jun 11 16:14:03 2020
In reply to: 15391
Reply to this: 15393
|
Author: |
gautam |
Type: |
Update |
Category: |
VAC |
Subject: |
Vac failure - probable cause is serial comm glitch |
|
|
Summary:
It looks like the main vacuum interlock was tripped due to a serial communication error from the TP2 controller. With Rana/Koji's permission, I will open V1 and expose the main volume to TP1 again (#2 in last section).
Details:
- The vacuum interlock log file at /opt/target/vac.log on c1vac suggests that the interlock was tripped because "TP2 is too warm".
- Looking back at the diagnostics channels, it looks like the TP2 temperature channel registered a rise in temperature of >30 C in <0.2 seconds, see Attachment #1 - seems highly unlikely, probably some kind of glitch in the serial communication? This particular pump is relatively new from Agilent (<2 years installed I think)
- The PSL shutter was automatically closed at ~1150 am today, see Attachment #2. There is some EPICS logic on c1psl (Acromag server) that checks if C1:Vac-P1a_pressure is greater than 3 mTorr (or greater than 500 Torr for in-air locking of the IMC), in which case it closes the shutter, so this seems consistent with expectations.
Recommended course of action:
- Code in some averaging in the interlock code, so that the interlock isn't triggered on some unphysical glitch like this. As shown in Attachment #3, this has been happening for the past 24 hours (though not before, because the interlock wasn't tripped). Probably need the derivative of the temperature as well, and the derivative should be less than 5 C/s or something physical (in addition to the temperature being high) for the interlock to trip.
- Re-open V1 to pump down the main volume to nominal pressure so that the interferometer locking activity can resume.
- One option in the interim is to bypass the TP2 temperature interlock condition.
- The pressure-based interlocks are probably sufficient to protect the main volume / pumps during the nominal operations - the temperature interlocks are mainly useful during the pumpdown where the TPs have a large load, and so we want to avoid over-stressing them.
|
|
|
|
|